I have a parallelogram (rv lot)45 feet by 30 feet with angle of 60 degrees..question is...will my rv of 40 feet by 10 feet fit on lot without crossing boundries?

Found 2 solutions by vleith, stanbon:
Answer by vleith(2983)

I think that's a trick question. Sure it will ... if you park it at a 60 degree angle.
If you have to park it 'straight', then we have to do a little math.
Draw your parallelogram, 45 feet and 30 feet with 60 degree angle between the 30 and 45 sides.
Drop a line from one end of the a 45 foot side down. Make it perpendicular to the 30 foot base line.
Now take 45 *cos(60) to see how long the side of the resulting right triangle is. Take 45*sin(60) to see how long the other side is. Then make the call about whether the rv can fit

Answer by stanbon(75887)

I have a parallelogram (rv lot)45 feet by 30 feet with angle of 60 degrees..question is...will my rv of 40 feet by 10 feet fit on lot without crossing boundries?
-------------------------
Draw the picture of the rv in the parallelogram.
The back of the rv is the base of an equilateral
triangle with 60 degree angles and sides of 10 ft.
------------
The triangle with the 120 degree angle has sides
of 20 ft and 45 ft.
Determine whether the 3rd side is >=40 ft. by
using the Law of Cosines
-----------
x^2 = 20^2 + 45^2 -2*20*45*cos120 = 3325
x = 57.66 ft.
-----------------
Yes your 40 foot long rv will fit in the lot.
=================
Cheers,
Stan H.